# A Unified Framework for Model Editing

### v1

@@ 3 Preservation-Memorization : A Unifying Framework for ROME and MEMIT
This is almost the same formula as equation 4 with minor changes. Here, $\Delta^{l}$ represents the change in weights at layer $l$, where $l\in\{L-(n-1),L-(n-2),\dots L\}$ represents one of the $n$ layers being edited. $V^{L}_{E}=V_{E}$ are the representations of the fact being edited at the final edit layer, which is represented by $L$. All other representations of $K_{E}$ and $C_{0}$ are calculated at the layer $l$ being edited. For $n=1$, the formula reduces to equation 4. We call this algorithm the edit-distribution algorithm, which is applied post-hoc after finding the closed-form solutions to the PM-objective.
In this paper, we want to disentangle these two different parts of MEMIT. The edit-distribution algorithm is separate from the solutions of the ROME and MEMIT objectives, therefore, we can apply the edit-distribution algorithm when using ROME, as well as use MEMIT without distributing the edits into multiple layers. The formula for using the MEMIT edit-distribution algorithm on ROME is as follows:
@@
To the best of our knowledge, none of the works so far differentiate between the MEMIT-objective and the edit-distribution algorithm, and as a consequence we never see edits using ROME being distributed to multiple layers or MEMIT being used on only a single layer. The unified code for the two algorithms can be found here [^4] https://github.com/scalable-model-editing/unified-model-editing. With this code-base, the MEMIT edit-distribution algorithm can be applied to ROME as well as EMMET which we introduce later.
3.1 Single Layer Edits Using ROME and MEMIT
In this section, we compare the ROME objective with the MEMIT objective, disentangling the effect of the edit-distribution algorithm. We make only single layer edits using the ROME and MEMIT objectives, without distributing the edits into multiple layers. Since ROME does not allow batched-edits, we evaluate models when only singular edits are made to them with both objectives. We evaluate three different models (GPT2-XL, GPT-J and Llama2-7B) on 1k examples selected at random from the CounterFact (Meng et al., 2022a) dataset. The implementation details about selected layers and hyperparameters are provided in section A.1.
The results are shown in Table 1. We see that solutions to both ROME and MEMIT objectives perform equally well at making singular edits across different metrics, without needing to distribute the edits to multiple layers. The opposite test to this would be to make multi-layer edits using both ROME and MEMIT. Edit-distribution algorithms are usually applied to improve performance of model editing when similar performance cannot be reached by just editing one layer. Since the edit scores are close to perfect for singular edits, applying edit-distribution algorithms are not useful and produce similar results. This can be seen in Table 2.
@@ 3.2 Impact of edit-distribution Algorithms
In the previous section, we see that both ROME and MEMIT objectives perform equally well when making singular edits to the model. We don’t see the usefulness of the edit-distribution algorithms when making singular edits. But MEMIT gives us the ability of making batched edits. In this section, we evaluate the performance of making batched edits using MEMIT as a function of the batch size on the CounterFact dataset. We find that as we increase batch-size, editing just a single layer in the model is less effective. Thus we need to edit multiple layers inside a model and hence need edit-distribution algorithms.
The results are shown in Figure 3. When only editing a single layer, we see that MEMIT is able to successfully make batched edits up to a batch size of 1024 for GPT2-XL, 256 for Llama-2-7b and a batch-size as large as 4096 for GPT-J. Here we define edit success by efficacy score, which measures if a new fact was successfully added or corrected inside the model. For GPT2-XL, we see that the edit score improves for batch size greater than 1024 with edit-distribution algorithms. The same is true for GPT-J beyond a batch size of 4096. If we consider all metrics together, as done in the "Score" metric, we can see improvements with edit-distribution starting at batch size 256 for GPT2-XL and GPT-J. This shows that edit-distribution algorithms become important as the batch size of the edits increase.
Throughout our experiments, we’ve found that ROME and MEMIT perform exceedingly well on GPT-J. While GPT-J is supposed to represent a larger model in the experiments shown in Meng et al. (2022a), it may also be the case that it is an easier model to edit. Because of this, we also perform single-layer and multi-layer editing experiments on Llama-2-7b. The selection of layers for Llama-2-7b has been taken from prior works (Yao et al., 2023; Zhang et al., 2024). We see that edit-distribution using the algorithm proposed in Meng et al. (2022b) hurts the model editing performance for Llama-2-7b, as shown in Figure 3.
With this experiment, we want to highlight two things - firstly, that edit-distribution algorithms are separate entities from the optimization objectives in ROME and MEMIT, and can be removed or added based on need. Secondly, the edit-distribution algorithm proposed in MEMIT is far from perfect. As seen for Llama-2-7b, using the MEMIT edit-distribution algorithm leads to loss of performance against single layer editing. We see lots of papers on model editing algorithms but almost none on edit-distribution algorithms, which were a big factor in editing memories in large batch sizes for MEMIT. We hope that this experiment can promote further research into the edit-distribution algorithms.
4 Introducing EMMET
In the previous section, we show that ROME and MEMIT are both algorithms optimizing the preservation-memorization objective of model editing, where ROME does memorization using an equality constraint wherease MEMIT uses a least-square objective for memorization. The least-square objective allows MEMIT to have a closed form solution for batched-editing, whereas with ROME, we can only edit one fact at a time. Thus, we ask the question - can we perform batched-editing under an equality constraint for memorization?
@@ 4 Introducing EMMET
$\underset{\hat{W}}{\operatorname{argmin}}\hskip 4.0pt\underbrace{\left\|\hat{W}K_{0}-W_{0}K_{0}\right\|}_{\text{preservation}}\hskip 8.0pt\text{ such that }\hskip 8.0pt\underbrace{\hat{W}k^{e}_{i}=v^{e}_{i}\hskip 8.0pt\forall i\in[1,2\dots E]}_{\text{memorization}}$ (7)
As can be seen in the above equation, the preservation of representations happens in the first term whereas memorization of all the new facts are forced using an equality constrain in the second term. The above equation is solved using the lagrange-multipliers. The Lagrangian for the above equation for multiple equality constraints requires a summation of lagrange multipliers and can be written as:
$L(\hat{W},\lambda_{1},\dots,{\lambda}_{E})=\frac{1}{2}\hat{W}K_{0}K^{T}_{0}\hat{W}^{T}-\hat{W}K_{0}K^{T}_{0}W^{T}_{0}+\frac{1}{2}W_{0}K_{0}K^{T}_{0}W^{T}_{0}-\sum_{i=1}^{E}\lambda_{i}^{T}(\hat{W}k^{e}_{i}-v^{e}_{i})$ (8)
@@
$\hskip 8.0pt\hat{W}=W_{0}+\Lambda K_{E}^{T}C^{-1}_{0}\$ (11)
To find the unknown matrix of lagrange multipliers ($\Lambda$), we make use of the equality constraint in EMMET. We know that $\hat{W}k^{e}_{i}=v^{e}_{i}\hskip 8.0pt\forall i\in[1,2\dots E]$, or $\hat{W}K_{E}=V_{E}$, where $V_{E}=[v^{e}_{1}\hskip 4.0pt|v^{e}_{2}\hskip 4.0pt|\dots|\hskip 4.0ptv^{e}_{E}]$. We replace equation 11 into the equality constraint to get:
$\Lambda K_{E}^{T}C^{-1}_{0}K_{E}+W_{0}K_{E}=V_{E}$ (12)
which can be re-arranged to get:
$\hskip 8.0pt\hat{W}=W_{0}+\Delta\hskip 10.0pt\text{where}\hskip 10.0pt\Delta=\left(V_{E}-W_{0}K_{E}\right)\left(K_{E}^{T}C^{-1}_{0}K_{E}\right)^{-1}K_{E}^{T}C^{-1}_{0}\$ (14)
@@
Figure 5: Performance of EMMET when editing a single layer compared to distributing the edit over multiple layers using the MEMIT edit-distribution algorithm on the CounterFact dataset.
We write the update equation of EMMET in a familiar form, where the residual $R=V_{E}-W_{0}K_{E}$ is modified by some matrix operations to update the models with new edits. Additionally, when we put $E=1$, the $K_{E}$ matrix reduces to a single vector $k_{e}$ and equation 14 reduces to the ROME update equation (equation 2). With EMMET, we complete the unification of ROME and MEMIT under the preservation-memorization objective and achieve a symmetry with the usage of these algorithms. EMMET allows for making batched-edits as well as singular when using equality constraints for memorization, much similar to MEMIT with least-square based memorization. The invertibility of matrix $B=K_{E}^{T}C^{-1}_{0}K_{E}$ cannot be guaranteed just like $C_{0}$, but we find that in practice, $B$ is usually invertible for smaller batch sizes like $C_{0}$.
@@ 4.1 Batch Editing with EMMET
We begin by experimenting with EMMET for model editing with varied batch sizes on GPT2-XL, GPT-J and Llama-2-7b on the CounterFact dataset. As shown previously, we disentangle the optimization objective with edit-distribution algorithms and show the comparison between EMMET and MEMIT when editing a single layer. We use identical configurations for both models including the layers being edited. The single layer editing comparison between EMMET and MEMIT can be found in Figure 4.
We find that EMMET performs competitvely with MEMIT up to a batch size of 256, after which the performance for EMMET begins to decline. We see a similar trend when we compare multi-layer experiments between EMMET and MEMIT using the MEMIT edit-distribution algorithm. The results can be seen in Figure 6. We suspect that this happens because the equality constraint in EMMET is too strong to edit a large number of (possibly conflicting) facts at the same time. While this shows that equality constraints for memorization are less effective for model editing for large batch sizes, increasing batch sizes is not the only way to scale model editing. EMMET can also be used for sequential editing (Yao et al., 2023; Gupta et al., 2024; Gupta and Anumanchipalli, 2024) with smaller batch sizes that fall within its competency range and provide a viable alternate for large scale model editing. We release EMMET to the research community and invite researchers to use it for large-scale model editing. The code base for EMMET is included in this repository - https://github.com/scalable-model-editing/unified-model-editing.
Finally, we compare the performance of EMMET when editing multiple layers using the edit-distribution algorithm proposed in MEMIT. The results can be seen in Figure 5. We make similar observations as before. For GPT2-XL, the edit-distribution algorithm seems to help improve editing performance. For GPT-J, single layer editing performs as well as multi-layer editing as it is an easier model to edit. For Llama-2-7b, we find that using the edit-distribution algorithm can hurt the model performance. This emphasizes our point of viewing edit-distribution algorithms as separate enities from the optimization objectives.
@@ 5 Conclusion
In this paper we present a conceptual framework that unites two popular model editing techniques - ROME and MEMIT, under the same objective. We call this objective the preservation-memorization objective, where we preserve certain representations of the model while memorizing the representations of a fact we want to store in a model. We show that ROME performs memorization using an equality-constraint, whereas MEMIT does so with a least-square constraint. We hope that this unifying framework improves the intuitive understanding of these algorithms and fuels future research based on both intuition and mathematics.
We also disentangle the edit-distribution algorithm proposed in MEMIT from the optimization objective and present them as separate entities. We show that while the MEMIT edit-distribution algorithm helps in editing performance for GPT2-XL and GPT-J at scale, using the same algorithm can also hurt performance as seen for Llama-2-7b. We thus encourage future research to view these as distinct entities and hope this promotes further research into edit-distribution algorithms. A fair comparison of future model editing techniques with MEMIT should be based on the objective of MEMIT and not conflating it with the edit-distribution algorithm.
Finally, we present EMMET - Equality-constrained Mass Model Editing in a Transformer. EMMET is a new batched-editing algorithm based on the preservation-memorization objective where batched-memorization happens under an equality constraint. Our experiments show that EMMET performs competitively with MEMIT for batch sizes up to 256. EMMET is unable to scale beyond this without loss of performance likely due to the hard equality constraints in its objective. Yet EMMET provides a viable alternative to batch editing at smaller batch sizes and can be used in combination with sequential editing to scale model editing.
@@ Appendix A Appendix
A.1 Implementation Details for ROME, MEMIT and EMMET
We use the standard implementation of ROME and MEMIT based on Meng et al. (2022a) and Meng et al. (2022b). The range of layers edited for GPT2-XL is $[13,17]$, for GPT-J is $[3-8]$ and for Llama-2-7b is $[4-8]$, with the final layer in the range being selected for single layer experiments. These choices are directly taken from Meng et al. (2022a) and Meng et al. (2022b) for GPT2-XL and GPT-J. We follow the work of Yao et al. (2023) for choices of layers and hyperparameters for llama-2-7b.

### v2

@@ 4 Edit-Distribution Algorithms
where $\Delta^{l}$ represents the change in weights at layer $l$, where $l\in\{L-(n-1),L-(n-2),\dots L\}$ represents one of the $n$ layers being edited. $V^{L}_{E}=V_{E}$ are the representations of the fact being edited at the final edit layer, which is represented by $L$. All other representations of $K_{E}$ and $C_{0}$ are calculated at the layer $l$ being edited. For $n=1$, the formula reduces to equation 8. We call this algorithm a type of edit-distribution algorithm, which is applied post-hoc after finding the closed-form solutions to the PM-objective.
The edit-distribution algorithm is separate from the solutions of the ROME and MEMIT objectives, therefore, we can apply the edit-distribution algorithm when using ROME, as well as use MEMIT without distributing the edits into multiple layers. The formula for using the MEMIT edit-distribution algorithm on ROME is as follows:
@@
Prior works on model editing do not differentiate between the MEMIT-objective and the edit-distribution algorithm, and as a consequence we never see edits using ROME being distributed to multiple layers or MEMIT being used on only a single layer.
@@ 4.1 Impact of edit-distribution Algorithms
The key advantage of the edit-distribution algorithm is apparent when making batched edits. In this section, we perform two experiments to analyze this. First, we compare single edits in ROME and MEMIT with and without edit distribution on 1k randomly selected facts from the CounterFact dataset. Following that, we compare batched editing in MEMIT with and without edit distribution. Both experiments are performed on three different models - GPT2-XL (1.5B), GPT-J (6B) and Llama2-7B.
The results are shown in Table 1 for edits without distribution and Table 3 for edits with distribution. We use the more stable version of ROME called r-ROME as presented in (Gupta and Anumanchipalli, 2024) that does not lead to model collapse and improves generalization. We see that solutions to both ROME and MEMIT objectives perform equally well at making singular edits across different metrics, without needing to distribute the edits to multiple layers. To highlight the usefulness of edit-distribution algorithms, we compare MEMIT when making batched edits. The results are shown in Figure 3. When only editing a single layer, we see that MEMIT is able to successfully make batched edits up to a batch size of 1024 for GPT2-XL, 256 for Llama-2-7b and a batch-size as large as 4096 for GPT-J [^3] In our experiments we find GPT-J to be an easier model to edit compared to other models. This is both intriguing but also not the best model to evaluate model editing success.. After this point, the performance of model editing increases when making edits on multiple layers, except for Llama-2-7b. All hyperparameters for all models were chosen as is from prior work (Meng et al., 2022a, b; Yao et al., 2023; Zhang et al., 2024) (appendix A.2).
With these experiments, we want to highlight two key points - firstly, when comparing the effectiveness of two optimization objectives, the evaluation should not be conflated with the edit distribution algorithms. Secondly, the MEMIT edit-distribution algorithm is not perfect and currently is the only way to distribute edits into multiple layers, where the residual in the update is distributed with specific ratios through different layers. We hope these experiments will bring more focus to edit distribution algorithms and boost further research in these methods.
@@
5 Introducing EMMET
In section 3, we show that ROME and MEMIT are both algorithms optimizing the preservation-memorization objective of model editing, where ROME does memorization using an equality constraint wherease MEMIT uses a least-square objective for memorization. Thus, we ask the question - can we perform batched-editing under an equality constraint for memorization?
@@ 5 Introducing EMMET
[…]
$\hskip 8.0pt\underbrace{\hat{W}k^{e}_{i}=v^{e}_{i}\hskip 8.0pt\forall i\in[1,2\dots E]}_{\text{memorization}}$
As can be seen in the above equation, the preservation of representations happens in the first term whereas memorization of all the new facts are forced using an equality constrain in the second term. The above equation is solved using lagrange-multipliers. The Lagrangian for the above equation for multiple equality constraints requires a summation of lagrange multipliers and equals:
$L(\hat{W},\lambda_{i})=\frac{1}{2}\hat{W}K_{0}K^{T}_{0}\hat{W}^{T}-\hat{W}K_{0}K^{T}_{0}W^{T}_{0}$ (12)
$+\frac{1}{2}W_{0}K_{0}K^{T}_{0}W^{T}_{0}-\sum_{i=1}^{E}\lambda_{i}^{T}(\hat{W}k^{e}_{i}-v^{e}_{i})$
@@
Figure 5: Performance comparison of EMMET and MEMIT when distributing the edit over multiple layers using the MEMIT edit-distribution algorithm on the CounterFact dataset.
$\hskip 8.0pt\hat{W}=W_{0}+\Lambda K_{E}^{T}C^{-1}_{0}\$ (15)
where $\Lambda=[\lambda_{1}\hskip 4.0pt|\lambda_{2}\hskip 4.0pt|\dots|\hskip 4.0pt\lambda_{E}]$, $K_{E}=[k^{e}_{1}\hskip 4.0pt|k^{e}_{2}\hskip 4.0pt|\dots|\hskip 4.0ptk^{e}_{E}]$ and $C_{0}=K_{0}K_{0}^{T}$.
The unknown matrix of lagrange multipliers ($\Lambda$) can be found using the constraint $\hat{W}K_{E}=V_{E}$ in the previous equation. It comes out to be:
$\hat{W}=W_{0}+\Delta\hskip 10.0pt\text{where}$ (17)
$\Delta$ $=\left(V_{E}-W_{0}K_{E}\right)\left(K_{E}^{T}C^{-1}_{0}K_{E}\right)^{-1}K_{E}^{T}C^{-1}_{0}$
We write the update equation of EMMET in a familiar form, where the residual $R=V_{E}-W_{0}K_{E}$ is modified by some matrix operations to update the models with new edits. Additionally, when we put $E=1$, the $K_{E}$ matrix reduces to a single vector $k_{e}$ and equation 17 reduces to the ROME update equation (equation 5). With EMMET, we complete the unification of ROME and MEMIT under the preservation-memorization objective and achieve a symmetry with the usage of these algorithms. EMMET allows for making batched-edits as well as singular when using equality constraints for memorization, much similar to MEMIT with least-square based memorization.
5.1 Stabilizing EMMET
There are two important matrices that are being inverted in EMMET and MEMIT. The first one is $C_{0}=K_{0}K^{T}_{0}$, which is defined identically in both algorithms, whereas $D=K_{E}^{T}C^{-1}_{0}K_{E}$ is only inverted in EMMET. While the invertibility of both matrices are assumed, they are not always guaranteed. Each of the matrices $K_{0}$ or $K_{E}$ can be written as a row of column vectors as explained in section 3, and thus $C_{0}$ can be written as a sum of outer products:
$C_{0}=K_{0}K^{T}_{0}=\sum_{i}k^{0}_{i}k^{0^{T}}_{i}$ (18)
where $k^{0}_{i}$ represent a key-vector we want to preserve. For an LLM of dimension $d$, the dimensionality of a key-vector is usually $4d$ (Figure 2), which is the dimensionality of the square matrix $C_{0}$. If $C_{0}$ is a $4d$-dimensional square matrix which is a summation of rank-1 matrices, it is invertible as long as there are atleast $4d$-independent vectors in the summation, or $4d$-independent vectors in $K_{0}$. For example, for Llama-2-7b with hidden dimension of 4096, the dimensionality of key vectors are 16384. So as long as representations of atleast 16384 independent key-vectors are being preserved while editing, $C_{0}$ will be an invertible matrix. In practice, we preserve representations of a much larger number of vectors, and hence this condition is almost always satisfied.
The matrix $D=K_{E}^{T}C^{-1}_{0}K_{E}$ is a square matrix of dimensionality equal to the number of edits. If given that $C_{0}$ is invertible, $D$ is invertible as long as $K_{E}$ is full-rank, which means all key-vectors corresponding to facts being memorized are independent of each other. While this is not guaranteed, it can be verified before editing and facts corresponding to non-independent keys can be removed from a batch. In practice, we do not find invertibility of D being an issue. However, we find that $D$ is often ill-conditioned, which means that the ratio of the largest and smallest eigenvalues of $D$ explodes. This doesn’t necessarily mean that the matrix is singular (non-invertible), but it does mean that numerical computations involving the matrix inverse are unstable and can lead to large numerical errors. To counter this, we set $D=D+\alpha I$, where $\alpha$ is set to 0.1 after an ablation over multiple batch sizes. This allows for stable batched edits using EMMET.
@@ 5.2 Batch Editing with EMMET
We begin by experimenting with EMMET for model editing with varied batch sizes on GPT2-XL, GPT-J and Llama-2-7b on the CounterFact and zsRE Levy et al. (2017) datasets. The exact implementation details can be found in section A.2. We compare the performance of EMMET and MEMIT on batch sizes up to 10,000 while editing both single (to directly compare the optimization objectives) and multiple layers. The single layer editing comparison between EMMET and MEMIT can be found in Figure 4. We see that both methods have almost identical performance in practice across different metrics. MEMIT performs slightly better than EMMET for Llama-2-7b, as indicated by ES, PS and S metrics. We then apply the MEMIT edit-distribution on EMMET and compare it with MEMIT. The results are shown in Figure 5. We see that in this case, EMMET performs slightly better than MEMIT for Llama-2-7b. The results on the zsRE dataset tell a similar story and can be seen in Figure 7 and 8. These results present EMMET as a viable new batched-editing algorithm.
(a) EMMET
(b) MEMIT
Figure 6: Downstream performance of post-edit Llama2-7b model for EMMET and MEMIT on four GLUE tasks. Batch index 0 refers to downstream performance before editing, with the performance of 5 independent edits of batch size 256.
Previous work (Gu et al., 2024; Gupta et al., 2024) has shown that model editing is often accompanied by model degradation. Gupta et al. (2024) show this by evaluating the edited model on downstream tasks from the popular GLUE benchmark Wang et al. (2018). We adopt their evaluation setting and evaluate both EMMET and MEMIT on four downstream tasks - sentiment analysis (SST2) Socher et al. (2013), paraphrase detection (MRPC) Dolan and Brockett (2005), natural language inference (NLI) (Dagan et al., 2005; Haim et al., 2006; Giampiccolo et al., 2007; Bentivogli et al., 2009) and linguistic acceptability classification Warstadt et al. (2019) for doing downstream evaluation. The results are shown in Figure 6 for a batch size of 256. The results for other batch sizes can be found in Appendix A.2. We find that both EMMET and MEMIT also degrade the model similarly.
Although EMMET is unable outperform MEMIT, it is an important piece in unifying model editing under the preservation-memorization framework. Both algorithms are able to make successful batched edits upto a batch-size of 10,000 and lead to similar model degradation. EMMET imposes a "theoretically" stronger memorization constraint, yet we do not see an improvement in editing efficacy. This indicates that we may be reaching the limit of model editing capabilities under the preservation-memorization objective.
@@ 6 Conclusion
In this paper we unite two popular model editing techniques, ROME and MEMIT, under the preservation-memorization objective, with ROME performing equality-constrained edits and MEMIT operating under a least-square constraint. We disentangle the edit-distribution algorithm proposed in MEMIT from the optimization objective, presenting them as separate entities, emphasizing that a fair comparison of future model editing techniques with MEMIT should be based on the objective of MEMIT rather than conflating it with the edit-distribution algorithm.
Finally, we present EMMET - Equality-constrained Mass Model Editing in a Transformer, a new batched-editing algorithm based on the preservation-memorization objective where batched-memorization happens under an equality constraint. Our experiments show that EMMET performs similarly to MEMIT across multiple dimensions and metrics. EMMET completes batched editing using both types of objectives and truly unifies model editing under the preservation memorization framework. We hope that this unifying framework improves the intuitive understanding of these algorithms and fuels future research based on both intuition and mathematics.
7 Limitations
While our technique may streamline error correction processes, it does not address deeper structural limitations within models, such as edited models inadvertently amplifying existing errors or introducing new inaccuracies. Furthermore, the effectiveness of our method varies depending on the complexity of the model architecture and the nature of the edited knowledge as evidenced by our experiments. Despite having a theoretically ‘stronger’ memorization objective, EMMET is not able to outperform MEMIT, which also indicates that we might have reached a saturation point for model editing using naive implementations of the preservation-memorization objective, underscoring the fact that significant progress is yet to be made in understanding edit distribution and its implications. Thus, while our work contributes to a deeper understanding of model behavior, it is essential to recognize and account for these limitations in the interpretation and application of our findings.
8 Ethical Considerations
While our model editing method allows users to effectively correct for errors or update facts in models, caution is warranted. Our technique also introduces concerns for potential misuse such as malicious actors inserting harmful or false knowledge in LLMs that is absent from the original training data. As such, we warn readers that LLMs should not be considered reliable knowledge bases.
@@ Appendix A Appendix
Batch Size Num Batches Total Edits
4 25 100
16 10 160
64 5 320
256 5 1280
1024 3 3072
4096 2 8192
10,000 1 10,000
Table 2: Statistics for batch size and number of batches used to create the numbers for this paper.
A.1 Related Work
Model editing methods can be broadly classified into two types - methods that add information in-context (Mitchell et al., 2022; Zhong et al., 2023; Cohen et al., 2023), and methods that modify the parameters of underlying model (De Cao et al., 2021; Mitchell et al., 2021; Meng et al., 2022a, b; Tan et al., 2023). Various model editing techniques have been proposed in the past that tackle this problem in different ways. Dai et al. (2021) first identify knowledge containing neurons in a model using integrated gradients (Sundararajan et al., 2017) and then modify the selected neurons to edit facts in a model. This method is not scalable with increasing model sizes as it requires us to find activations for each neuron in the model. De Cao et al. (2021) and Mitchell et al. (2021) train a hypernetwork (Chauhan et al., 2023) that generates the new weights of the model being edited. While these methods have been optimized to scale with a square-root dependence on the size of the edited model, it still requires training of additional editing models dependent on each source model being edited. Other methods add the most relevant updated knowledge in context (Mitchell et al., 2022; Cohen et al., 2023; Zhong et al., 2023). While such methods provide a viable alternative to model editing, in this paper, we focus on parameter-modifying model editing methods, namely ROME Meng et al. (2022a) and Meng et al. (2022b).
(a) Efficacy Accuracy (EM)
(b) Paraphrase Accuracy (PM)
(c) Neighborhood Accuracy (NM)
Figure 7: Single layer editing performance of EMMET as a function of batch size when compared to MEMIT on the zsRE dataset.